# On an unconditional $\rm GL_{3}$ analog of Selberg’s result

### v1

@@ 3. An approximation of $S_{F}(t)$
$\begin{split}J_{1}=&-\int_{\sigma_{x}}^{\infty}\operatorname{Im}\frac{L^{\prime}}{L}(\sigma+it,F)\mathrm{d}\sigma\\ =&\int_{\sigma_{x}}^{\infty}\operatorname{Im}\sum_{n\leq x^{3}}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{\sigma+it}}\mathrm{d}\sigma+O\left(\left|\sum_{n\leq x^{3}}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{\sigma_{x}+it}}\right|\int_{\sigma_{x}}^{\infty}x^{1/4-\sigma/2}\mathrm{d}\sigma\right)\\ &+O\left(\log(|t|+M_{F})\int_{\sigma_{x}}^{\infty}x^{1/4-\sigma/2}\mathrm{d}\sigma\right)\\ =&\operatorname{Im}\sum_{n\leq x^{3}}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{\sigma_{x}+it}\log n}+O\left(\frac{1}{\log x}\left|\sum_{n\leq x^{3}}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{\sigma_{x}+it}}\right|\right)+O\left(\frac{\log(|t|+M_{F})}{\log x}\right).\end{split}$
@@
$\begin{split}J_{2}\leq&\,(\sigma_{x}-1/2)\left|\frac{L^{\prime}}{L}(\sigma_{x}+it,F)\right|\\ \ll&\,(\sigma_{x}-1/2)\left|\sum_{n\leq x^{3}}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{\sigma_{x}+it}}\right|+(\sigma_{x}-1/2)\log(|t|+M_{F}).\end{split}$
@@
$\begin{split}&\operatorname{Im}\left(\frac{L^{\prime}}{L}(\sigma_{x}+it,F)-\frac{L^{\prime}}{L}(\sigma+it,F)\right)\\ =&\sum_{\rho}\frac{(\sigma_{x}-\sigma)(\sigma+\sigma_{x}-2\beta)(t-\gamma)}{\big{(}(\sigma_{x}-\beta)^{2}+(t-\gamma)^{2}\big{)}\big{(}(\sigma-\beta)^{2}+(t-\gamma)^{2}\big{)}}+O(1).\end{split}$
@@
$\begin{split}\int_{1/2}^{\sigma_{x}}\frac{|\sigma+\sigma_{x}-2\beta||t-\gamma|}{(\sigma-\beta)^{2}+(t-\gamma)^{2}}\mathrm{d}\sigma&\leq 3(\sigma_{x}-1/2)\int_{-\infty}^{\infty}\frac{|t-\gamma|}{(\sigma-\beta)^{2}+(t-\gamma)^{2}}\mathrm{d}\sigma\\ &\leq 10(\sigma_{x}-1/2).\end{split}$ (3.15)
@@
$\begin{split}\int_{1/2}^{\sigma_{x}}\frac{|\sigma+\sigma_{x}-2\beta||t-\gamma|}{(\sigma-\beta)^{2}+(t-\gamma)^{2}}\mathrm{d}\sigma&<\int_{1/2}^{\sigma_{x}}\frac{|\sigma+\sigma_{x}-2\beta|}{|t-\gamma|}\mathrm{d}\sigma\\ &\leq\int_{1/2}^{\sigma_{x}}\frac{6|\beta-1/2|}{3|\beta-1/2|}\mathrm{d}\sigma=2(\sigma_{x}-1/2).\end{split}$ (3.16)
By the bounds (3.15)–(3.16) and Lemma 3.3, we obtain
$\begin{split}|J_{3}|\leq&10(\sigma_{x}-1/2)\sum_{\rho}\frac{\sigma_{x}-1/2}{(\sigma_{x}-\beta)^{2}+(t-\gamma)^{2}}+O(\sigma_{x}-1/2)\\ =&O\left((\sigma_{x}-1/2)\left|\sum_{n\leq x^{3}}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{\sigma_{x}+it}}\right|\right)+O\big{(}(\sigma_{x}-1/2)\log(|t|+M_{F})\big{)}.\end{split}$
Thus we complete the proof of lemma. ∎
@@
Let $t>0$ be fixed and $T$ be any large parameter. For $x=T^{\delta/3}$ with $0<\delta<\frac{3\theta}{8nk+3}$, $k\in\mathbb{N}$, we have
@@
$\begin{split}&(\sigma_{x,F}-1/2)^{4n}x^{4nk(\sigma_{x,F}-1/2)}\\ \leq&\left(\frac{10}{\log x}\right)^{4n}x^{40nk/\log x}+2^{4n+1}\sum_{\beta>\frac{1}{2}+\frac{5}{\log x}\atop|t-\gamma|\leq\frac{x^{3|\beta-1/2|}}{\log x}}(\beta-1/2)^{4n}x^{8nk(\beta-1/2)}.\end{split}$
@@
$\begin{split}&\sum_{\beta>\frac{1}{2}+\frac{5}{\log x}\atop|t-\gamma|\leq\frac{x^{3|\beta-1/2|}}{\log x}}(\beta-1/2)^{4n}x^{8nk(\beta-1/2)}\\ \leq&\sum_{i=5}^{\frac{1}{2}\lfloor\log x\rfloor}\left(\frac{i+1}{\log x}\right)^{4n}x^{8nk\frac{i+1}{\log x}}\sum_{\frac{1}{2}+\frac{i}{\log x}<\beta\leq\frac{1}{2}+\frac{i+1}{\log x}\atop|t-\gamma|\leq\frac{x^{3|\beta-1/2|}}{\log x}}1\\ \leq&\frac{1}{(\log x)^{4n}}\sum_{i=5}^{\frac{1}{2}\lfloor\log x\rfloor}\left(i+1\right)^{4n}e^{8nk(i+1)}N\left(\frac{i}{\log x},|t|+\frac{e^{3(i+1)}}{\log x};F\right).\end{split}$
@@
$\begin{split}&\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}\sum_{\beta>\frac{1}{2}+\frac{5}{\log x}\atop|t-\gamma|\leq\frac{x^{3|\beta-1/2|}}{\log x}}(\beta-1/2)^{4n}x^{8nk(\beta-1/2)}\\ \leq&\,\frac{1}{(\log x)^{4n}}\sum_{i=5}^{\frac{1}{2}\lfloor\log x\rfloor}\left(i+1\right)^{4n}e^{8nk(i+1)}\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}N\left(\frac{i}{\log x},|t|+\frac{e^{3(i+1)}}{\log x};F\right)\\ \ll&\,\frac{1}{(\log x)^{4n}}\sum_{i=5}^{\infty}\left(i+1\right)^{4n}e^{8nk(i+1)}\left(|t|+\frac{e^{3(i+1)}}{\log x}\right)T^{-\theta\frac{i}{\log x}}\log T\\ \ll&{}_{t,n,\delta,k}\,\frac{\log T}{(\log x)^{4n+1}}\sum_{i=5}^{\infty}\left(i+1\right)^{4n}e^{(8nk+3-\frac{3\theta}{\delta})i}\\ \ll&{}_{t,n,\delta,k}\,\frac{1}{(\log T)^{4n}}\end{split}$
provided that $0<\delta<\frac{3\theta}{8nk+3}$. In addition, by Corollary 2.3,
$\begin{split}\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}\left(\frac{10}{\log x}\right)^{4n}x^{40nk/\log x}\ll_{n,k}\frac{1}{(\log x)^{4n}}\ll_{n,k,\delta}\frac{1}{(\log T)^{4n}}.\end{split}$
Thus we complete the proof of lemma. ∎
@@
$\mathcal{M}_{F}(t):=\frac{1}{\pi}\operatorname{Im}\sum_{p\leq x^{3}}\frac{C_{F}(p)}{p^{1/2+it}}\qquad\text{and}\qquad\mathcal{R}_{F}(t):=S_{F}(t)-\mathcal{M}_{F}(t),$ (4.1)
@@
$\begin{split}\mathcal{R}_{F}(t)=&O\left(\left|\operatorname{Im}\sum_{p\leq x^{3}}\frac{A_{F}(1,p)(\Lambda_{x}(p)-\Lambda(p))}{p^{1/2+it}\log p}\right|\right)+O\left(\sum_{k=2}^{7}\left|\operatorname{Im}\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})}{p^{k(1/2+it)}\log p}\right|\right)\\ &+O\left((\sigma_{x}-1/2)\sum_{k=1}^{7}x^{k(\sigma_{x}-1/2)}\int_{1/2}^{\infty}x^{k(1/2-\sigma)}\left|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+it)}}\right|\mathrm{d}\sigma\right)\\ &+O\big{(}(\sigma_{x}-1/2)\log(|t|+M_{F})\big{)}+O\big{(}1\big{)}.\end{split}$
@@
$\begin{split}\mathcal{R}_{F}(t)=&\frac{1}{\pi}\operatorname{Im}\sum_{p\leq x^{3}}\frac{C_{F}(p)(\Lambda_{x}(p)p^{1/2-\sigma_{x}}-\Lambda(p))}{p^{1/2+it}\log p}+\frac{1}{\pi}\operatorname{Im}\sum_{k=2}^{\infty}\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})}{p^{k(\sigma_{x}+it)}\log p^{k}}\\ &+O\left((\sigma_{x}-1/2)\left|\sum_{k=1}^{\infty}\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})}{p^{k(\sigma_{x}+it)}}\right|\right)+O\big{(}(\sigma_{x}-1/2)\log(|t|+M_{F})\big{)}.\end{split}$
@@
$\begin{split}\sum_{k=8}^{\infty}\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})}{p^{k(\sigma_{x}+it)}}=O(1),\quad\text{and}\quad\sum_{k=8}^{\infty}\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})}{p^{k(\sigma_{x}+it)}\log p}=O(1).\end{split}$
@@
$\begin{split}\mathcal{R}_{F}(t)=&\,\frac{1}{\pi}\operatorname{Im}\sum_{p\leq x^{3}}\frac{A_{F}(1,p)(\Lambda_{x}(p)-\Lambda(p))}{p^{1/2+it}\log p}+\frac{1}{\pi}\operatorname{Im}\sum_{k=2}^{7}\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})}{p^{k(1/2+it)}\log p^{k}}\\ &+\frac{1}{\pi}\operatorname{Im}\sum_{k=1}^{7}\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})(p^{k(1/2-\sigma_{x})}-1)}{p^{k(1/2+it)}\log p^{k}}\\ &+O\left((\sigma_{x}-1/2)\left|\sum_{k=1}^{7}\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})}{p^{k(\sigma_{x}+it)}}\right|\right)\\ &+O\big{(}(\sigma_{x}-1/2)\log(|t|+M_{F})\big{)}+O(1).\end{split}$
@@
$\begin{split}\left|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})}{p^{k(a+it)}}\right|&=kx^{k(a-1/2)}\left|\int_{a}^{\infty}x^{k(1/2-\sigma)}\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+it)}}\mathrm{d}\sigma\right|\\ &\leq kx^{k(\sigma_{x}-1/2)}\int_{1/2}^{\infty}x^{k(1/2-\sigma)}\left|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+it)}}\right|\mathrm{d}\sigma.\end{split}$
@@
$\begin{split}&\left|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})(p^{k(1/2-\sigma_{x})}-1)}{p^{k(1/2+it)}\log p^{k}}\right|=\left|\int_{1/2}^{\sigma_{x}}\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})}{p^{k(a+it)}}\mathrm{d}a\right|\\ &\leq k(\sigma_{x}-1/2)x^{k(\sigma_{x}-1/2)}\int_{1/2}^{\infty}x^{k(1/2-\sigma)}\left|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+it)}}\right|\mathrm{d}\sigma,\end{split}$
@@
$\begin{split}\left|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})}{p^{k(\sigma_{x}+it)}}\right|\leq kx^{k(\sigma_{x}-1/2)}\int_{1/2}^{\infty}x^{k(1/2-\sigma)}\left|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+it)}}\right|\mathrm{d}\sigma.\end{split}$
Thus we complete the proof of lemma. ∎
@@
Let $t>0$ be fixed and $T$ be any large parameter. For $x=T^{\delta/3}$ with sufficiently small $\delta>0$, we have
@@
$\mathcal{M}_{F}(t)=\frac{1}{\pi}\operatorname{Im}\sum_{p\leq x^{3}}\frac{C_{F}(p)}{p^{1/2+it}}=\frac{-i}{2\pi}\left(\sum_{p\leq x^{3}}\frac{A_{F}(1,p)}{p^{1/2+it}}-\sum_{p\leq x^{3}}\frac{\overline{A_{F}(1,p)}}{p^{1/2-it}}\right).$
@@
$\mathcal{M}_{F}(t)^{n}=\frac{(-i)^{n}}{(2\pi)^{n}}\left(\sum_{p\leq T^{\delta}}\frac{A_{F}(1,p)}{p^{1/2+it}}-\sum_{p\leq T^{\delta}}\frac{\overline{A_{F}(1,p)}}{p^{1/2-it}}\right)^{n}$
@@
$\frac{A_{F}(1,p_{1})^{m(p_{1})}\overline{A_{F}(1,p_{1})^{n(p_{1})}}}{p_{1}^{m(p_{1})(1/2+it)}(-1)^{n(p_{1})}p_{1}^{n(p_{1})(1/2-it)}}\times\cdots\times\frac{A_{F}(1,p_{r})^{m(p_{r})}\overline{A_{F}(1,p_{r})^{n(p_{r})}}}{p_{r}^{m(p_{r})(1/2+it)}(-1)^{n(p_{r})}p_{r}^{n(p_{r})(1/2-it)}},$ (4.4)
@@
[…]

### v2

@@ 3. An approximation of $S_{F}(t)$
$\begin{split}J_{1}=&-\int_{\sigma_{x}}^{\infty}\operatorname{Im}\frac{L^{\prime}}{L}(\sigma+\mathrm{i}t,F)\mathrm{d}\sigma\\ =&\int_{\sigma_{x}}^{\infty}\operatorname{Im}\sum_{n\leq x^{3}}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{\sigma+\mathrm{i}t}}\mathrm{d}\sigma+O\bigg(\bigg|\sum_{n\leq x^{3}}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{\sigma_{x}+\mathrm{i}t}}\bigg|\int_{\sigma_{x}}^{\infty}x^{1/4-\sigma/2}\mathrm{d}\sigma\bigg)\\ &+O\bigg(\log(|t|+M_{F})\int_{\sigma_{x}}^{\infty}x^{1/4-\sigma/2}\mathrm{d}\sigma\bigg)\\ =&\,\operatorname{Im}\sum_{n\leq x^{3}}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{\sigma_{x}+\mathrm{i}t}\log n}+O\bigg(\frac{1}{\log x}\bigg|\sum_{n\leq x^{3}}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{\sigma_{x}+\mathrm{i}t}}\bigg|\bigg)+O\bigg(\frac{\log(|t|+M_{F})}{\log x}\bigg).\end{split}$
@@
$\begin{split}J_{2}\leq&\,(\sigma_{x}-1/2)\bigg|\frac{L^{\prime}}{L}(\sigma_{x}+\mathrm{i}t,F)\bigg|\\ \ll&\,(\sigma_{x}-1/2)\bigg|\sum_{n\leq x^{3}}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{\sigma_{x}+\mathrm{i}t}}\bigg|+(\sigma_{x}-1/2)\log(|t|+M_{F}).\end{split}$
@@
$\begin{split}&\operatorname{Im}\bigg(\frac{L^{\prime}}{L}(\sigma_{x}+\mathrm{i}t,F)-\frac{L^{\prime}}{L}(\sigma+\mathrm{i}t,F)\bigg)\\ =&\sum_{\rho}\frac{(\sigma_{x}-\sigma)(\sigma+\sigma_{x}-2\beta)(t-\gamma)}{\big((\sigma_{x}-\beta)^{2}+(t-\gamma)^{2}\big)\big((\sigma-\beta)^{2}+(t-\gamma)^{2}\big)}+O(1).\end{split}$
@@
$\int_{1/2}^{\sigma_{x}}\frac{|\sigma+\sigma_{x}-2\beta||t-\gamma|}{(\sigma-\beta)^{2}+(t-\gamma)^{2}}\mathrm{d}\sigma$ $\leq 3(\sigma_{x}-1/2)\int_{-\infty}^{\infty}\frac{|t-\gamma|}{(\sigma-\beta)^{2}+(t-\gamma)^{2}}\mathrm{d}\sigma$
$\leq 10(\sigma_{x}-1/2).$ (3.15)
@@
$\int_{1/2}^{\sigma_{x}}\frac{|\sigma+\sigma_{x}-2\beta||t-\gamma|}{(\sigma-\beta)^{2}+(t-\gamma)^{2}}\mathrm{d}\sigma$ $<\int_{1/2}^{\sigma_{x}}\frac{|\sigma+\sigma_{x}-2\beta|}{|t-\gamma|}\mathrm{d}\sigma$
$\leq\int_{1/2}^{\sigma_{x}}\frac{6|\beta-1/2|}{3|\beta-1/2|}\mathrm{d}\sigma=2(\sigma_{x}-1/2).$ (3.16)
By the estimates in (3)–(3) and Lemma 3.3, we obtain
$\begin{split}|J_{3}|\leq&10(\sigma_{x}-1/2)\sum_{\rho}\frac{\sigma_{x}-1/2}{(\sigma_{x}-\beta)^{2}+(t-\gamma)^{2}}+O(\sigma_{x}-1/2)\\ =&O\bigg((\sigma_{x}-1/2)\bigg|\sum_{n\leq x^{3}}\frac{C_{F}(n)\Lambda_{x}(n)}{n^{\sigma_{x}+\mathrm{i}t}}\bigg|\bigg)+O\big((\sigma_{x}-1/2)\log(|t|+M_{F})\big).\end{split}$
Thus we complete the proof of the theorem. ∎
@@
Let $t>0$ be fixed and $T$ be any large positive parameter. For $k\in\mathbb{N}$, $x=T^{\delta/3}$ with $0<\delta<\frac{3\theta}{8nk+3}$, we have
@@
$\begin{split}&(\sigma_{x,F}-1/2)^{4n}x^{4nk(\sigma_{x,F}-1/2)}\\ \leq&\bigg(\frac{10}{\log x}\bigg)^{4n}x^{40nk/\log x}+2^{4n+1}\sum_{\beta>\frac{1}{2}+\frac{5}{\log x}\atop|t-\gamma|\leq\frac{x^{3|\beta-1/2|}}{\log x}}(\beta-1/2)^{4n}x^{8nk(\beta-1/2)}.\end{split}$
@@
$\begin{split}&\sum_{\beta>\frac{1}{2}+\frac{5}{\log x}\atop|t-\gamma|\leq\frac{x^{3|\beta-1/2|}}{\log x}}(\beta-1/2)^{4n}x^{8nk(\beta-1/2)}\\ \leq&\sum_{i=5}^{\frac{1}{2}\lfloor\log x\rfloor}\bigg(\frac{i+1}{\log x}\bigg)^{4n}x^{8nk\frac{i+1}{\log x}}\sum_{\frac{1}{2}+\frac{i}{\log x}<\beta\leq\frac{1}{2}+\frac{i+1}{\log x}\atop|t-\gamma|\leq\frac{x^{3|\beta-1/2|}}{\log x}}1\\ \leq&\frac{1}{(\log x)^{4n}}\sum_{i=5}^{\frac{1}{2}\lfloor\log x\rfloor}\big(i+1\big)^{4n}\mathrm{e}^{8nk(i+1)}N\bigg(\frac{i}{\log x},|t|+\frac{\mathrm{e}^{3(i+1)}}{\log x};F\bigg).\end{split}$
@@
$\begin{split}&\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}\sum_{\beta>\frac{1}{2}+\frac{5}{\log x}\atop|t-\gamma|\leq\frac{x^{3|\beta-1/2|}}{\log x}}(\beta-1/2)^{4n}x^{8nk(\beta-1/2)}\\ \leq&\,\frac{1}{(\log x)^{4n}}\sum_{i=5}^{\frac{1}{2}\lfloor\log x\rfloor}\big(i+1\big)^{4n}\mathrm{e}^{8nk(i+1)}\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}N\bigg(\frac{i}{\log x},|t|+\frac{\mathrm{e}^{3(i+1)}}{\log x};F\bigg)\\ \ll&\,\frac{1}{(\log x)^{4n}}\sum_{i=5}^{\infty}\big(i+1\big)^{4n}\mathrm{e}^{8nk(i+1)}\bigg(|t|+\frac{\mathrm{e}^{3(i+1)}}{\log x}\bigg)T^{-\theta\frac{i}{\log x}}\log T\\ \ll&{}_{t,n,\delta,k}\,\frac{\log T}{(\log x)^{4n+1}}\sum_{i=5}^{\infty}\big(i+1\big)^{4n}\mathrm{e}^{(8nk+3-\frac{3\theta}{\delta})i}\\ \ll&{}_{t,n,\delta,k}\,\frac{1}{(\log T)^{4n}}\end{split}$
provided that $0<\delta<\frac{3\theta}{8nk+3}$. In addition, by Corollary 2.3 and (1.9),
$\begin{split}\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}\bigg(\frac{10}{\log x}\bigg)^{4n}x^{40nk/\log x}\ll_{n,k}\frac{1}{(\log x)^{4n}}\ll_{n,k,\delta}\frac{1}{(\log T)^{4n}}.\end{split}$
Thus we complete the proof of the lemma. ∎
@@
$\mathcal{M}_{F}(t):=\frac{1}{\pi}\operatorname{Im}\sum_{p\leq x^{3}}\frac{C_{F}(p)}{p^{1/2+\mathrm{i}t}}\qquad\text{and}\qquad\mathcal{R}_{F}(t):=S_{F}(t)-\mathcal{M}_{F}(t),$ (4.1)
@@
$\begin{split}\mathcal{R}_{F}(t)=&O\bigg(\bigg|\operatorname{Im}\sum_{p\leq x^{3}}\frac{A_{F}(1,p)(\Lambda_{x}(p)-\Lambda(p))}{p^{1/2+\mathrm{i}t}\log p}\bigg|\bigg)+O\bigg(\sum_{k=2}^{7}\bigg|\operatorname{Im}\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})}{p^{k(1/2+\mathrm{i}t)}\log p}\bigg|\bigg)\\ &+O\bigg((\sigma_{x}-1/2)\sum_{k=1}^{7}x^{k(\sigma_{x}-1/2)}\int_{1/2}^{\infty}x^{k(1/2-\sigma)}\bigg|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+\mathrm{i}t)}}\bigg|\mathrm{d}\sigma\bigg)\\ &+O\big((\sigma_{x}-1/2)\log(|t|+M_{F})\big)+O\big(1\big).\end{split}$
@@
$\begin{split}\mathcal{R}_{F}(t)=&\frac{1}{\pi}\operatorname{Im}\sum_{p\leq x^{3}}\frac{C_{F}(p)(\Lambda_{x}(p)p^{1/2-\sigma_{x}}-\Lambda(p))}{p^{1/2+\mathrm{i}t}\log p}+\frac{1}{\pi}\operatorname{Im}\sum_{k=2}^{\infty}\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})}{p^{k(\sigma_{x}+\mathrm{i}t)}\log p^{k}}\\ &+O\bigg((\sigma_{x}-1/2)\bigg|\sum_{k=1}^{\infty}\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})}{p^{k(\sigma_{x}+\mathrm{i}t)}}\bigg|\bigg)+O\big((\sigma_{x}-1/2)\log(|t|+M_{F})\big).\end{split}$
@@
$\begin{split}\sum_{k=8}^{\infty}\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})}{p^{k(\sigma_{x}+\mathrm{i}t)}}=O(1),\quad\text{and}\quad\sum_{k=8}^{\infty}\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})}{p^{k(\sigma_{x}+\mathrm{i}t)}\log p}=O(1).\end{split}$
@@
$\begin{split}\mathcal{R}_{F}(t)=&\,\frac{1}{\pi}\operatorname{Im}\sum_{p\leq x^{3}}\frac{A_{F}(1,p)(\Lambda_{x}(p)-\Lambda(p))}{p^{1/2+\mathrm{i}t}\log p}+\frac{1}{\pi}\operatorname{Im}\sum_{k=2}^{7}\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})}{p^{k(1/2+\mathrm{i}t)}\log p^{k}}\\ &+\frac{1}{\pi}\operatorname{Im}\sum_{k=1}^{7}\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})(p^{k(1/2-\sigma_{x})}-1)}{p^{k(1/2+\mathrm{i}t)}\log p^{k}}\\ &+O\bigg((\sigma_{x}-1/2)\bigg|\sum_{k=1}^{7}\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})}{p^{k(\sigma_{x}+\mathrm{i}t)}}\bigg|\bigg)\\ &+O\big((\sigma_{x}-1/2)\log(|t|+M_{F})\big)+O(1).\end{split}$
@@
$\begin{split}\bigg|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})}{p^{k(a+\mathrm{i}t)}}\bigg|&=kx^{k(a-1/2)}\bigg|\int_{a}^{\infty}x^{k(1/2-\sigma)}\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+\mathrm{i}t)}}\mathrm{d}\sigma\bigg|\\ &\leq kx^{k(\sigma_{x}-1/2)}\int_{1/2}^{\infty}x^{k(1/2-\sigma)}\bigg|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+\mathrm{i}t)}}\bigg|\mathrm{d}\sigma.\end{split}$
@@
$\begin{split}&\bigg|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})(p^{k(1/2-\sigma_{x})}-1)}{p^{k(1/2+\mathrm{i}t)}\log p^{k}}\bigg|=\bigg|\int_{1/2}^{\sigma_{x}}\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})}{p^{k(a+\mathrm{i}t)}}\mathrm{d}a\bigg|\\ &\leq k(\sigma_{x}-1/2)x^{k(\sigma_{x}-1/2)}\int_{1/2}^{\infty}x^{k(1/2-\sigma)}\bigg|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+\mathrm{i}t)}}\bigg|\mathrm{d}\sigma,\end{split}$
@@
$\begin{split}\bigg|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})}{p^{k(\sigma_{x}+\mathrm{i}t)}}\bigg|\leq kx^{k(\sigma_{x}-1/2)}\int_{1/2}^{\infty}x^{k(1/2-\sigma)}\bigg|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+\mathrm{i}t)}}\bigg|\mathrm{d}\sigma.\end{split}$
Thus we complete the proof of the lemma. ∎
@@
Let $t>0$ be fixed and $T$ be any large positive parameter. For $x=T^{\delta/3}$ with sufficiently small $\delta>0$, we have
@@
$\mathcal{M}_{F}(t)=\frac{1}{\pi}\operatorname{Im}\sum_{p\leq x^{3}}\frac{C_{F}(p)}{p^{1/2+\mathrm{i}t}}=\frac{-\mathrm{i}}{2\pi}\bigg(\sum_{p\leq x^{3}}\frac{A_{F}(1,p)}{p^{1/2+\mathrm{i}t}}-\sum_{p\leq x^{3}}\frac{\overline{A_{F}(1,p)}}{p^{1/2-\mathrm{i}t}}\bigg).$
@@
$\mathcal{M}_{F}(t)^{n}=\frac{(-\mathrm{i})^{n}}{(2\pi)^{n}}\bigg(\sum_{p\leq T^{\delta}}\frac{A_{F}(1,p)}{p^{1/2+\mathrm{i}t}}-\sum_{p\leq T^{\delta}}\frac{\overline{A_{F}(1,p)}}{p^{1/2-\mathrm{i}t}}\bigg)^{n}$
@@
$\frac{A_{F}(1,p_{1})^{m(p_{1})}\overline{A_{F}(1,p_{1})^{n(p_{1})}}}{p_{1}^{m(p_{1})(1/2+\mathrm{i}t)}(-1)^{n(p_{1})}p_{1}^{n(p_{1})(1/2-\mathrm{i}t)}}\times\cdots\times\frac{A_{F}(1,p_{r})^{m(p_{r})}\overline{A_{F}(1,p_{r})^{n(p_{r})}}}{p_{r}^{m(p_{r})(1/2+\mathrm{i}t)}(-1)^{n(p_{r})}p_{r}^{n(p_{r})(1/2-\mathrm{i}t)}},$ (4.4)
@@
Case (I): In the general term (4.4), $m(p_{j_{0}})\not\equiv n(p_{j_{0}})(\bmod\,3)$ for some $j_{0}$. By Lemma 2.1, we have